Prove That:
(tan^2 A + 1)Cos2A= 2-Sec^2A

Answer by mananth(16946)


(tan^2 A + 1)Cos2A= 2-Sec^2A

(tan^2A+1)(cos^2A-sin^2A)

sec^2A *(cos^2A-(1-cos^2A)

sec^2A *(cos^2A-1+cos^2A)

1/cos^2A *2cos^2A -1/cos^2A

=2-sec^2A